The area of a rectangular parking lot os 1800 sq m. The width is 14 m less than the length. What are the dimensions?
Answer by mananth(16949)

Let the length of the plot be x
the width will be x-14
Area of plot= L * W
x(x-14)= 1800
x^2-14x=1800
x^2-14x-1800=0
x^2-50x+36x-1800=0
x(x-50)+36(x-50)=0
(x-50)(x+36)=0
x=50 which is the length
50-14 = 36 is the width